solve: (2x+3)3 +(x-1)3=2(3x-2)2+3(15x2+5x+34)
Answer by palanisamy(496)

(2x+3)3 +(x-1)3=2(3x-2)2+3(15x2+5x+34)
6x+9+3x-3 = 12x-8+90+15x+102
9x+6 = 27x+184
9x-27x = 184-6
-18x = 178
x = (178)/-18
x = -89/9